The lifetime of a cellphone is normally distributed with a mean of 18 months and a standard deviation of 2.8 months. What is the probability a cell phone lasts less than 10.88 months? 
Answer #1

M=18, 8.2.1, x= 10.88 pcx< 10.88) = P(ze wody =P(ze 10.56-10) = prz<2.5429 - 2.8 - Joooss [options]

7 .00 .01 .02 -3.4 .0003 ,0003 ,0003 -3.3 .0005 .0005 .0005 -3.2 .0007 .0007 .0006 -3.1 .0010 .0009 .0009 -3.0 .0013 .0013 .0